Grade 11Algebra

THE TOTAL number of ways in which 4 persons can be chosen from 16 persons sitting in a circle such that no two of the choosen are neighbours..... Pnc

Nitin Bindal
9 Years agoGrade 11

1 Answer

Saurabh Koranglekar
6 Years ago
The solution is for total 10 people and instead of 4 persons 3 are chosen

First find the total number of ways of selecting three people out of ten seated in a round table. Total number of selection is equal to10C3.
But the condition is that no two of them should be adjacent.
So first find the selections in which exactly two people are next to each other. Two people next to each other can be selected in10ways(AB,BC,CD,DE,EF,FG,GH,HI,IJ,JA). Once an adjacent pair of people selected, the remaining one person can be selected in 6 ways (so that he/she is not next to any of them).
So, total number of selections possible so that exactly two people are next to each other is equal to10×6=60
Also number ofselections in whichthree people are together is10.
(ABC, BCD........)
Hence, selections in which no two people are next to each other=total selection−( selections in which exactly two are together+selections in which all three are together)
=10C3​−(60+10)=120−70=50